# Neural Discrete Token Representation Learning for Extreme Token Reduction in Video Large Language Models

### v4

@@ 1 Introduction
We attribute these challenges to three key limitations. First, existing methods rely on fixed-count or fixed-percentage reduction strategies, which either leave sequences overly long, with redundant tokens, or prune so aggressively that critical information is lost. Second, they lack adaptive, context-sensitive mechanisms for selecting the most informative tokens in the frames. Third, none leverage vector quantization to cluster tokens into discrete categories, hindering substantial gains in information density through thorough compression. To address these limitations, we propose VQToken, a vector-quantized token representation framework that dynamically clusters continuous ViT embeddings into a compact, discrete codebook. By mapping each token to its nearest codebook entry, VQToken produces a minimal set of discrete tokens while preserving spatial–temporal relationships. Accurately capturing spatial–temporal dynamics within this discrete clustering, however, remains a critical challenge.
The second major challenge is preserving spatial–temporal coherence during token reduction. Traditional pruning methods Kim et al. (2022); Liu et al. (2024b) often discard positional cues that are vital for tracking object motion accurately. Likewise, similarity-based merging techniques Bolya et al. (2023); Lee et al. (2024); Li et al. (2024a) tend to ignore spatial–temporal encodings or reapply them inconsistently, which undermines dynamic context modeling. To overcome this, we introduce a token-hashing mechanism based on vector quantization. First, each grid-level token is mapped to its nearest codebook centroid; then, we record its original $(f,h,w)$ index in a three-dimensional hash table. This table integrates seamlessly into the VQToken architecture, preserving positional information in a compact discrete form. By leveraging the inherent redundancy of video data. Inspired by computationally expensive motion tracking techniques like optical flow, our token-hashing mechanism offers a lightweight alternative that preserves essential spatial–temporal context with minimal computational overhead.
@@
    A formal definition of the Extreme Token Reduction task, together with the Token Information Density (TokDense) metric and separate complexity measures for the reduction module and downstream LLM inference, covering both fixed-length and adaptive-length settings.
@@
    Empirical evidence that VQToken compresses video token sequences to just 0.07% of their original length with only a 0.66% drop in NextQA-MC accuracy, achieving leading efficiency and information density while maintaining competitive performance across multiple benchmarks.
@@ 2.2 Token Reduction
Figure 1: Overview of neural discrete token representation learning. First, the input video is tokenized into a continuous sequence of visual tokens. An adaptive discrete process then clusters and vector-quantizes these tokens into a compact codebook. A token hash function records each token’s original spatial–temporal location and maps it to its nearest codebook entry. The VQ-Attention module integrates the codebook with the index map to produce a compressed token sequence that preserves positional information. Finally, the compressed tokens and a tokenized query are passed to the large language model for zero-shot inference.
@@ 3.1 Problem Definition: Extreme Token Reduction
TokDense=AccuracyToken $\mathrm{TokDense}=\frac{\mathrm{Accuracy}}{\text{Token Count}},$ (2)
@@ 3.2 Neural Discrete Token Representation Learning
We introduce Neural Discrete Token Representation Learning, a vector-quantization architecture that dynamically minimizes token sequence length while preserving complete spatial–temporal motion information.
@@ 3.2.1 Adaptive Discrete Process
Unlike fixed-ratio merging methods such as ToMe Bolya et al. (2023), which risk under-merging or spurious groupings, our adaptive discrete process selects the number of clusters either statically or dynamically. For fixed-length reduction, we use classical K-Means Vassilvitskii and Arthur (2006); for adaptive-length reduction, we employ an adaptive K-Means variant Bhatia et al. (2004). While video-segmentation approaches (e.g., SAM-based Ravi et al. (2024)) can yield fine-grained clusters, their computational overhead makes them less practical for this stage.
@@
3.2.2 Vector-Quantization Architecture
To transform the discrete clusters into a compact token sequence while preserving spatial–temporal information, we design three components: a concise codebook, a token hash function, and a VQ-based reduction module.
@@ 3.2.2 Vector-Quantization Architecture
Token-Hash Fuction Mapping.
@@
VQ-Based Reduction Module.
We integrate the codebook $B$ and index map $M$ via a lightweight VQ-Attention mechanism using a lightweight VQ-Attention block that enriches each centroid with motion context without increasing token count,:
@@ 4.1 Implementation Details
We follow the LLaVA-OneVision Li et al. (2024a) setup, using 178K videos paired with 1.3M instruction-following samples (178K captions, 960K open-ended questions, 196K multiple-choice questions) to cover diverse video scenarios.
@@
We report Accuracy (Acc.), the percentage of correct responses on multiple-choice and open-ended QA tasks; Token Count, the number of tokens processed per example; Throughput, measured in frames per second; FLOPs (T), the total tera-FLOPs required for inference; and Run-Time, the end-to-end inference latency. To disentangle the cost of token reduction from its downstream benefit, we also measure Module Complexity, the time complexity of the reduction module alone, and LLM Complexity, the complexity of the large language model given the reduced token sequence. Finally, Token Information Density (TokDense)—defined in Eq. 2—quantifies accuracy per retained token.
@@
Due to limited computational resources, we evaluate on the 0.5B-parameter track, using LLaVA-OV-SI and LLaVA-OneVision as baselines. For efficiency, we integrate our VQToken framework with LLaVA-OneVision to minimize GPU usage. We also compare against 7B-parameter versions; despite having 14× more parameters and greater compute, our 0.5B model—using only 0.14% of the original token count—outperforms some 7B models, highlighting our method’s efficiency.
@@
For token reduction, we compare against several baselines: Token Pruning Kim et al. (2022): A widely recognized method for reducing token numbers and increasing throughput in LLMs. Token Merging (ToMe) Bolya et al. (2023): A popular baseline for token reduction, known for its efficiency improvements. Video Token Merging Lee et al. (2024): The current state-of-the-art method for token reduction in video large language models, extending the capabilities of ToMe to video data. Interpolation: Introduced by Li et al. (2024a), the use of bilinear interpolation to reduce the number of tokens in visual representations, particularly for video frames. This approach allows the model to handle a larger number of frames by reducing the tokens per frame, achieving a balance between performance and computational cost. (v) DyCoke Tao et al. (2024), the current SOTA method that employs temporal compression to merge redundant tokens across frames and dynamic KV-cache reduction to selectively prune spatial redundancy.
@@ 4.3 Extreme Token Reduction Task
To evaluate efficiency under extreme compression, we compare VQ-Token against classical and state-of-the-art reduction methods—Token Pruning, ToMe, and VidToMe—using fixed token budgets. We configure our model with a predetermined cluster size $K$ and set each baseline to retain exactly the same number of tokens (e.g., 12, 32, or 64). This controlled setting isolates the effect of each reduction strategy on accuracy. As Table 3 shows, VQ-Token consistently outperforms frame-level merging (ToMe), sequence-level merging (VidToMe), and pruning across all extreme budgets, demonstrating superior preservation of semantic content under severe token constraints.
@@
Here, each method dynamically selects the optimal token count based on video content complexity. We report both accuracy and the average tokens used per sample (“Avg Tokens”). Table 3 illustrates that, compared to interpolation-based downsampling and our own fixed-length variant, the adaptive VQ-Token model achieves higher accuracy while consuming significantly fewer tokens on average. This result underscores its ability to balance efficiency and performance in a content-aware manner.
@@ 4.5 Performance on Multiple Benchmarks
Model #Parameters Year Zero-Shot Acc.(%) $\uparrow$ Token Num.$\downarrow$
@@ 4.6 Ablation Study
We evaluate each component’s contribution by incrementally adding the discrete codebook, token hash function (Indices), and VQ attention to the LLaVA-OV baseline. As shown in Table 6, the Base alone compresses tokens to 0.14% but incurs a 22.0% accuracy drop, highlighting the loss of spatial–temporal cues. Incorporating the Indices marginally improves accuracy, although the LLM cannot yet leverage motion information effectively. Introducing Attn restores accuracy substantially, demonstrating that VQ attention is essential to integrate positional context into the compressed tokens. Randomizing each module’s parameters (denoted “rand”) leads to significant performance degradation, confirming the necessity of properly learned codebook, mapping, and attention for extreme token reduction.
@@
Figure 3 illustrates clustering behaviors on sample frames. We compare adaptive K-Means on token embeddings with the state-of-the-art mask-based segmentation model Segment Anything (SAM) Ravi et al. (2024). Both methods group semantically similar regions and maintain cluster consistency across frames. While SAM yields finer-grained regions, adaptive K-Means offers a more computationally efficient alternative that sufficiently captures object trajectories for the token hash function. This efficiency makes adaptive K-Means the preferred choice for vLLMs requiring extreme compression with minimal overhead.
@@ 5 Conclusion
We have introduced VQToken, the first neural discrete token representation framework to leverage adaptive vector quantization for extreme token reduction in video large language models. VQToken constructs a compact codebook from continuous ViT embeddings and preserves spatial–temporal positions via a hash-based token mapping, enabling plug-and-play integration with existing architectures. To benchmark extreme compression, we formalized the Extreme Token Reduction task and proposed the Token Information Density (TokDense) metric, along with separate complexity measures for the reduction module and downstream LLM inference. These contributions provide a comprehensive evaluation suite for both fixed-length and adaptive-length reduction settings. Empirically, VQToken compresses token sequences to just 0.07% of their original length, achieving over 99.9% reduction, with only a 0.66% drop in NextQA-MC accuracy. It matches comparable performance on ActivityNet-QA, Long Video Bench, and VideoMME, while delivering state-of-the-art efficiency and information density. Ablation studies confirm that the codebook, token hash function, and VQ-attention are all critical to preserving semantic and motion information under extreme compression. Efficiency analyses demonstrate substantial reductions in FLOPs, latency, token information density, and overall computational complexity compared to prior methods. In future work, we will explore hierarchical clustering and learned cluster-size schedules to further optimize compression, as well as extend the VQToken framework to downstream tasks such as video generation and motion prediction.

### v5

# VQToken: Neural Discrete Token Representation Learning for Extreme Token Reduction in Video Large Language Models

@@ 1 Introduction
We attribute these challenges to three key limitations. First, existing methods rely on fixed‐count or fixed‐percentage reduction strategies, which either leave sequences overly long, with redundant tokens, or prune so aggressively that critical information is lost. Second, they lack adaptive, context‐sensitive mechanisms for selecting the most informative tokens in the frames. Third, none leverage vector quantization to cluster tokens into discrete categories, hindering substantial gains in information density through thorough compression. To address these limitations, we propose VQToken, a vector‐quantized token representation framework that dynamically clusters continuous ViT embeddings into a compact, discrete codebook. By mapping each token to its nearest codebook entry, VQToken produces a minimal set of discrete tokens while preserving spatial–temporal relationships. Accurately capturing spatial–temporal dynamics within this discrete clustering, however, remains a critical challenge.
The second major challenge is preserving spatial–temporal coherence during token reduction. Traditional pruning methods Kim et al. (2022); Liu et al. (2024b) often discard positional cues that are vital for tracking object motion accurately. Likewise, similarity‐based merging techniques Bolya et al. (2023); Lee et al. (2024); Li et al. (2024a) tend to ignore spatial–temporal encodings or reapply them inconsistently, which undermines dynamic context modeling. To overcome this, we introduce a token‐hashing mechanism based on vector quantization. First, each grid‐level token is mapped to its nearest codebook centroid; then, we record its original $(f,h,w)$ index in a three‐dimensional hash table. This table integrates seamlessly into the VQToken architecture, preserving positional information in a compact discrete form. By leveraging the inherent redundancy of video data. Inspired by computationally expensive motion tracking techniques like optical flow, our token‐hashing mechanism offers a lightweight alternative that preserves essential spatial–temporal context with minimal computational overhead.
@@
    A formal definition of the Extreme Token Reduction task, together with the Token Information Density (TokDense) metric and separate complexity measures for the reduction module and downstream LLM inference, covering both fixed‐length and adaptive‐length settings.
@@
    Empirical evidence that VQToken compresses video token sequences to just 0.07% of their original length with only a 0.66% drop in NextQA‐MC accuracy, achieving leading efficiency and information density while maintaining competitive performance across multiple benchmarks.
@@ 2.2 Token Reduction
Figure 1: Overview of neural discrete token representation learning. First, the input video is tokenized into a continuous sequence of visual tokens. An adaptive discrete process then clusters and vector‐quantizes these tokens into a compact codebook. A token hash function records each token’s original spatial–temporal location and maps it to its nearest codebook entry. The VQ‐Attention module integrates the codebook with the index map to produce a compressed token sequence that preserves positional information. Finally, the compressed tokens and a tokenized query are passed to the large language model for zero‐shot inference.
@@ 3.1 Problem Definition: Extreme Token Reduction
$\mathrm{TokDense}=\frac{\mathrm{Accuracy}}{\text{Token Count}},$ (2)
@@ 3.2 Neural Discrete Token Representation Learning
We introduce Neural Discrete Token Representation Learning, a vector‐quantization architecture that dynamically minimizes token sequence length while preserving complete spatial–temporal motion information.
@@ 3.2.1 Adaptive Discrete Process
Unlike fixed‐ratio merging methods such as ToMe Bolya et al. (2023), which risk under‐merging or spurious groupings, our adaptive discrete process selects the number of clusters either statically or dynamically. For fixed‐length reduction, we use classical K‐Means Vassilvitskii and Arthur (2006); for adaptive‐length reduction, we employ an adaptive K‐Means variant Bhatia et al. (2004). While video‐segmentation approaches (e.g., SAM-based Ravi et al. (2024)) can yield fine-grained clusters, their computational overhead makes them less practical for this stage.
@@
3.2.2 Vector‐Quantization Architecture
To transform the discrete clusters into a compact token sequence while preserving spatial–temporal information, we design three components: a concise codebook, a token hash function, and a VQ‐based reduction module.
@@ 3.2.2 Vector‐Quantization Architecture
Token‐Hash Fuction Mapping.
@@
VQ‐Based Reduction Module.
We integrate the codebook $B$ and index map $M$ via a lightweight VQ‐Attention mechanism using a lightweight VQ‐Attention block that enriches each centroid with motion context without increasing token count,:
@@ 4.1 Implementation Details
We follow the LLaVA-OneVision Li et al. (2024a) setup, using 178K videos from LLaVA-Video-178K dataset paired with 1.3M instruction-following samples (178K captions, 960K open-ended questions, 196K multiple-choice questions) to cover diverse video scenarios.
@@
We report Accuracy (Acc.), the percentage of correct responses on multiple‐choice and open‐ended QA tasks; Token Count, the number of tokens processed per example; Throughput, measured in frames per second; FLOPs (T), the total tera‐FLOPs required for inference; and Run‐Time, the end‐to‐end inference latency. To disentangle the cost of token reduction from its downstream benefit, we also measure Module Complexity, the time complexity of the reduction module alone, and LLM Complexity, the complexity of the large language model given the reduced token sequence. Finally, Token Information Density (TokDense)—defined in Eq. 2—quantifies accuracy per retained token.
@@
Due to limited computational resources, we evaluate on the 0.5B‐parameter track, using LLaVA‐OV‐SI and LLaVA‐OneVision as baselines. For efficiency, we integrate our VQToken framework with LLaVA‐OneVision to minimize GPU usage. We also compare against 7B‐parameter versions; despite having 14× more parameters and greater compute, our 0.5B model—using only 0.14% of the original token count—outperforms some 7B models, highlighting our method’s efficiency.
@@
For token reduction, we compare against several baselines: Token Pruning Kim et al. (2022): A widely recognized method for reducing token numbers and increasing throughput in LLMs. Token Merging (ToMe) Bolya et al. (2023): A popular baseline for token reduction, known for its efficiency improvements. Video Token Merging Lee et al. (2024): The current state-of-the-art method for token reduction in video large language models, extending the capabilities of ToMe to video data. Interpolation: Introduced by Li et al. (2024a), the use of bilinear interpolation to reduce the number of tokens in visual representations, particularly for video frames. This approach allows the model to handle a larger number of frames by reducing the tokens per frame, achieving a balance between performance and computational cost. (v) DyCoke Tao et al. (2024), the current SOTA method that employs temporal compression to merge redundant tokens across frames and dynamic KV‐cache reduction to selectively prune spatial redundancy.
@@ 4.3 Extreme Token Reduction Task
To evaluate efficiency under extreme compression, we compare VQ-Token against classical and state‐of‐the‐art reduction methods—Token Pruning, ToMe, and VidToMe—using fixed token budgets. We configure our model with a predetermined cluster size $K$ and set each baseline to retain exactly the same number of tokens (e.g., 12, 32, or 64). This controlled setting isolates the effect of each reduction strategy on accuracy. As Table 3 shows, VQ-Token consistently outperforms frame‐level merging (ToMe), sequence‐level merging (VidToMe), and pruning across all extreme budgets, demonstrating superior preservation of semantic content under severe token constraints.
@@
Here, each method dynamically selects the optimal token count based on video content complexity. We report both accuracy and the average tokens used per sample (“Avg Tokens”). Table 3 illustrates that, compared to interpolation‐based downsampling and our own fixed‐length variant, the adaptive VQ-Token model achieves higher accuracy while consuming significantly fewer tokens on average. This result underscores its ability to balance efficiency and performance in a content‐aware manner.
@@ 4.5 Performance on Multiple Benchmarks
Model #Parameters Year Zero-Shot Acc.(%) $\uparrow$ Token Num.% $\downarrow$
@@ 4.6 Ablation Study
We evaluate each component’s contribution by incrementally adding the discrete codebook, token hash function (Indices), and VQ attention to the LLaVA‐OV baseline. As shown in Table 6, the Base alone compresses tokens to 0.14% but incurs a 22.0% accuracy drop, highlighting the loss of spatial–temporal cues. Incorporating the Indices marginally improves accuracy, although the LLM cannot yet leverage motion information effectively. Introducing Attn restores accuracy substantially, demonstrating that VQ attention is essential to integrate positional context into the compressed tokens. Randomizing each module’s parameters (denoted “rand”) leads to significant performance degradation, confirming the necessity of properly learned codebook, mapping, and attention for extreme token reduction.
@@
Figure 3 illustrates clustering behaviors on sample frames. We compare adaptive K‐Means on token embeddings with the state-of-the-art mask‐based segmentation model Segment Anything (SAM) Ravi et al. (2024). Both methods group semantically similar regions and maintain cluster consistency across frames. While SAM yields finer-grained regions, adaptive K‐Means offers a more computationally efficient alternative that sufficiently captures object trajectories for the token hash function. This efficiency makes adaptive K‐Means the preferred choice for vLLMs requiring extreme compression with minimal overhead.
@@ 5 Conclusion
We have introduced VQToken, the first neural discrete token representation framework to leverage adaptive vector quantization for extreme token reduction in video large language models. VQToken constructs a compact codebook from continuous ViT embeddings and preserves spatial–temporal positions via a hash‐based token mapping, enabling plug‐and‐play integration with existing architectures. To benchmark extreme compression, we formalized the Extreme Token Reduction task and proposed the Token Information Density (TokDense) metric, along with separate complexity measures for the reduction module and downstream LLM inference. These contributions provide a comprehensive evaluation suite for both fixed‐length and adaptive‐length reduction settings. Empirically, VQToken compresses token sequences to just 0.07% of their original length, achieving over 99.9% reduction, with only a 0.66% drop in NextQA-MC accuracy. It matches comparable performance on ActivityNet-QA, Long Video Bench, and VideoMME, while delivering state‐of‐the‐art efficiency and information density. Ablation studies confirm that the codebook, token hash function, and VQ‐attention are all critical to preserving semantic and motion information under extreme compression. Efficiency analyses demonstrate substantial reductions in FLOPs, latency, token information density, and overall computational complexity compared to prior methods. In future work, we will explore hierarchical clustering and learned cluster‐size schedules to further optimize compression, as well as extend the VQToken framework to downstream tasks such as video generation and motion prediction.